Suppose, in a two sector model comprising consumers and firms, that the individuals receive
the following payments for their provision of goods and services: wages K60000, interest
K30000, rent K10000 and profits k8000. Consumption spending is K101000, planned
investment expenditure is K20000 and actual investment expenditure is K7000.
a) What is the equilibrium output?
b) What is the level of saving in this economy; clearly show how it’s calculated.
c) What is the relationship of saving and investment?
d) What is a leakage from the circular flow of income?
e) Why is investment spending viewed as an injection into the circular flow of
income?
1
Expert's answer
2018-02-26T09:47:07-0500
Suppose, in a two sector model comprising consumers and firms, that the individuals receive
the following payments for their provision of goods and services: wages K60000, interest
K30000, rent K10000 and profits k8000. Consumption spending is K101000, planned
investment expenditure is K20000 and actual investment expenditure is K7000.
a) The equilibrium output is Y = C + I = 101000 + 7000 = 108000.
b) The level of saving in this economy is S = I = 7000.
c) Saving and investment should be equal, so S = I = 7000.
d) A leakage from the circular flow of income is the amount of saving.
e) Investment spending is viewed as an injection into the circular flow of
income, because it allows to increase income.